# A parallel framework for graphical optimal transport

### v1

@@ Appendix C Proofs of Lemmas
$-\log\left(\|q_{j}^{t+1}\|_{1}\right)\geq 1-\|q_{j}^{t+1}\|_{1}=\frac{1}{|N(j)|}\sum_{k\in N(j)}\|q_{(j,k)}^{t}\|_{1}-\|q_{j}^{t+1}\|_{1}$ (75)
$=$ $\mathbf{1}^{\top}\left(\frac{1}{|N(j)|}\sum_{k\in N(j)}q_{(j,k)}^{t}-q_{j}^{t+1}\right)\geq\frac{1}{11|N(j)|}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}^{2},$ (76)
where the last inequality is proved in (Kroshnin et al. 2019, proof of Lemma 6). Therefore,
$f(u^{t},\rho^{t})-f(u^{t+1},\rho^{t+1})$ $\geq\frac{1}{2}\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}^{2}+\frac{1}{11}\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}^{2}$ (77)
$\geq\frac{1}{11}\left(\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}^{2}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}^{2}\right)$ (78)
$\geq\frac{1}{22|E|}\left(\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}\right)^{2},$ (79)
where we use Cauchy-Schwarz in the final inequality. ∎
Proof.
Proof of Lemma 5 Denote the partition where the nodes have mismatch as $S$. Without loss of generality, we assume $j\in S$ and $k\in V\backslash S$.
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}=\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}-\widehat{B}_{(j,k)}{\rangle}$ (80)
$=\sum_{j\in\Gamma}{\langle}C_{(j,k_{j})},B_{(j,k_{j})}-\widehat{B}_{(j,k_{j})}{\rangle}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}{\langle}C_{(j,k)},B_{(j,k)}-\widehat{B}_{(j,k)}{\rangle}.$ (81)
@@
$\sum_{j\in\Gamma}{\langle}C_{(j,k_{j})},B_{(j,k_{j})}-\widehat{B}_{(j,k_{j})}{\rangle}\leq\sum_{j\in\Gamma}\|C_{(j,k_{j})}\|_{\infty}\|B_{(j,k_{j})}-\widehat{B}_{(j,k_{j})}\|_{1}\leq 2C_{\infty}\sum_{j\in\Gamma}\|\mu_{j}-B_{(j,k)}{\bf 1}\|_{1}.$ (82)
@@
$~{}~{}~{}~{}\sum_{j\notin\Gamma}\sum_{k\in N(j)}{\langle}C_{(j,k)},B_{(j,k)}-\widehat{B}_{(j,k)}{\rangle}\leq\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|C_{(j,k)}\|_{\infty}\|B_{(j,k)}-\widehat{B}_{(j,k)}\|_{1}$ (83)
$\leq 2\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|C_{(j,k)}\|_{\infty}\|\bar{q}_{j}-B_{(j,k)}{\bf 1}\|_{1}\leq 2C_{\infty}\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}-B_{(j,k)}{\bf 1}\|_{1}.$ (84)
@@
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}$ (85)
$\leq$ $2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-B_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(B)-B_{(j,k)}{\bf 1}\|_{1}\right).$ (86)
It follows the exact same way to obtain the same bound for $\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}$. ∎
Proof.
Proof of Lemma 6 Let $\{\widehat{Y}_{(j,k)}\}$ denote the output from Algorithm 4 with inputs $\{B^{*}_{(j,k)}\}$ and $\{\widetilde{B}_{(j,k_{j})}{\bf 1}\}_{j\in\Gamma}$. Note that $\widetilde{B}_{(j,k)}$ is the optimal solution to
$\min_{B\in\Pi(\widetilde{B}_{(j,k)}{\bf 1},~{}\widetilde{B}^{\top}_{(j,k)}{\bf 1})}{\langle}C_{(j,k)},B{\rangle}+\epsilon{\mathcal{H}}(B|M_{(j,k)}),\quad\text{for }\forall(j,k)\in E$ (87)
@@
${\langle}C_{(j,k)},\widetilde{B}_{(j,k)}{\rangle}+\epsilon{\mathcal{H}}(\widetilde{B}_{(j,k)}|M_{(j,k)})\leq{\langle}C_{(j,k)},\widehat{Y}_{(j,k)}{\rangle}+\epsilon{\mathcal{H}}(\widehat{Y}_{(j,k)}|M_{(j,k)}).$ (88)
@@
$~{}~{}~{}~{}{\langle}C_{(j,k)},\widetilde{B}_{(j,k)}{\rangle}-{\langle}C_{(j,k)},\widehat{Y}_{(j,k)}{\rangle}\leq\epsilon{\mathcal{H}}(\widehat{Y}_{(j,k)}|M_{(j,k)})-\epsilon{\mathcal{H}}(\widetilde{B}_{(j,k)}|M_{(j,k)})$ (89)
$\leq-\epsilon{\langle}\widetilde{B}_{(j,k)},\log(\widetilde{B}_{(j,k)}){\rangle}+\epsilon{\langle}\widetilde{B}_{(j,k)}-\widehat{Y}_{(j,k)},\log M_{(j,k)}{\rangle}$ (90)
$\leq 2\epsilon\log(d)+\epsilon{\langle}\log\gamma_{j},\widetilde{B}_{(j,k)}{\bf 1}-\widehat{Y}_{(j,k)}{\bf 1}{\rangle}+\epsilon{\langle}\log\gamma_{k},\widetilde{B}_{(j,k)}^{\top}{\bf 1}-\widehat{Y}_{(j,k)}^{\top}{\bf 1}{\rangle}=2\epsilon\log(d),$ (91)
@@
The plans $\{B^{*}_{(j,k)}\}$ already satisfy the condition $\bar{q}_{j}(B^{*})=B^{*}_{(j,k)}{\bf 1}$ for $\forall j\in\Gamma,k\in N(j)$. Thus Lemma 5 gives
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{Y}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B^{*}_{(j,k)}{\rangle}$ (92)
@@
$\leq 2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-\widetilde{B}_{(j,k_{j})}{\bf 1})\|_{1}\right),$ (93)
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widetilde{B}_{(j,k)}{\rangle}$ (94)
$\leq 2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-\widetilde{B}_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(\widetilde{B})-\widetilde{B}_{(j,k)}{\bf 1}\|_{1}\right).$ (95)
Summing up 91, 93 and 95 concludes the proof. ∎

### v2

@@ 9 Proofs of Lemmas
$-\log\left(\|q_{j}^{t+1}\|_{1}\right)\geq 1-\|q_{j}^{t+1}\|_{1}=\frac{1}{|N(j)|}\sum_{k\in N(j)}\|q_{(j,k)}^{t}\|_{1}-\|q_{j}^{t+1}\|_{1}$
$=$ $\mathbf{1}^{\top}\left(\frac{1}{|N(j)|}\sum_{k\in N(j)}q_{(j,k)}^{t}-q_{j}^{t+1}\right)\geq\frac{1}{11|N(j)|}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}^{2},$
where the last inequality is proved in [kroshnin2019complexity, proof of Lemma 6]. Therefore,
$f(u^{t},\rho^{t})-f(u^{t+1},\rho^{t+1})$ $\geq\frac{1}{2}\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}^{2}+\frac{1}{11}\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}^{2}$
$\geq\frac{1}{11}\left(\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}^{2}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}^{2}\right)$
$\geq\frac{1}{22|E|}\left(\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}\right)^{\!2}\!,$
where we use Cauchy-Schwarz in the final inequality.
Proof 9.5 (Proof of Lemma 3.9).
Denote the partition where the nodes have mismatch as $S$. Without loss of generality, we assume $j\in S$ and $k\in V\backslash S$.
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}=\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}-\widehat{B}_{(j,k)}{\rangle}$
$=\sum_{j\in\Gamma}{\langle}C_{(j,k_{j})},B_{(j,k_{j})}-\widehat{B}_{(j,k_{j})}{\rangle}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}{\langle}C_{(j,k)},B_{(j,k)}-\widehat{B}_{(j,k)}{\rangle}.$
@@
$\sum_{j\in\Gamma}{\langle}C_{(j,k_{j})},B_{(j,k_{j})}-\widehat{B}_{(j,k_{j})}{\rangle}$ $\leq\sum_{j\in\Gamma}\|C_{(j,k_{j})}\|_{\infty}\|B_{(j,k_{j})}-\widehat{B}_{(j,k_{j})}\|_{1}$
$\leq 2C_{\infty}\sum_{j\in\Gamma}\|\mu_{j}-B_{(j,k)}{\bf 1}\|_{1}.$
@@
$\sum_{j\notin\Gamma}\sum_{k\in N(j)}{\langle}C_{(j,k)},B_{(j,k)}-\widehat{B}_{(j,k)}{\rangle}\leq\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|C_{(j,k)}\|_{\infty}\|B_{(j,k)}-\widehat{B}_{(j,k)}\|_{1}$
$\leq 2\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|C_{(j,k)}\|_{\infty}\|\bar{q}_{j}-B_{(j,k)}{\bf 1}\|_{1}\leq 2C_{\infty}\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}-B_{(j,k)}{\bf 1}\|_{1}.$
@@
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}$
$\leq$ $2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-B_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(B)-B_{(j,k)}{\bf 1}\|_{1}\right).$
It follows the exact same way to obtain the same bound for $\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}$.
Proof 9.6 (Proof of Lemma 3.10).
Let $\{\widehat{Y}_{(j,k)}\}$ denote the output from Algorithm 3 with inputs $\{B^{*}_{(j,k)}\}$ and $\{\widetilde{B}_{(j,k_{j})}{\bf 1}\}_{j\in\Gamma}$. Note that $\widetilde{B}_{(j,k)}$ is the optimal solution to
$\min_{B\in\Pi(\widetilde{B}_{(j,k)}{\bf 1},~\widetilde{B}^{\top}_{(j,k)}{\bf 1})}{\langle}C_{(j,k)},B{\rangle}+\epsilon{\mathcal{H}}(B|M_{(j,k)}),\quad\text{for }\forall(j,k)\in E$
@@
${\langle}C_{(j,k)},\widetilde{B}_{(j,k)}{\rangle}+\epsilon{\mathcal{H}}(\widetilde{B}_{(j,k)}|M_{(j,k)})\leq{\langle}C_{(j,k)},\widehat{Y}_{(j,k)}{\rangle}+\epsilon{\mathcal{H}}(\widehat{Y}_{(j,k)}|M_{(j,k)}).$
@@
${\langle}C_{(j,k)},\widetilde{B}_{(j,k)}{\rangle}-{\langle}C_{(j,k)},\widehat{Y}_{(j,k)}{\rangle}\leq\epsilon{\mathcal{H}}(\widehat{Y}_{(j,k)}|M_{(j,k)})-\epsilon{\mathcal{H}}(\widetilde{B}_{(j,k)}|M_{(j,k)})$
$\leq-\epsilon{\langle}\widetilde{B}_{(j,k)},\log(\widetilde{B}_{(j,k)}){\rangle}+\epsilon{\langle}\widetilde{B}_{(j,k)}-\widehat{Y}_{(j,k)},\log M_{(j,k)}{\rangle}$
$\leq 2\epsilon\log(d)+\epsilon{\langle}\log\gamma_{j},\widetilde{B}_{(j,k)}{\bf 1}-\widehat{Y}_{(j,k)}{\bf 1}{\rangle}+\epsilon{\langle}\log\gamma_{k},\widetilde{B}_{(j,k)}^{\top}{\bf 1}-\widehat{Y}_{(j,k)}^{\top}{\bf 1}{\rangle}=2\epsilon\log(d),$
@@
The plans $\{B^{*}_{(j,k)}\}$ already satisfy the condition $\bar{q}_{j}(B^{*})=B^{*}_{(j,k)}{\bf 1}$ for $\forall j\in\Gamma,k\in N(j)$. Thus Lemma 3.9 gives
@@
(52) $\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{Y}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B^{*}_{(j,k)}{\rangle}\leq 2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-\widetilde{B}_{(j,k_{j})}{\bf 1})\|_{1}\right),$
(53) $\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widetilde{B}_{(j,k)}{\rangle}$
$\qquad\leq 2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-\widetilde{B}_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(\widetilde{B})-\widetilde{B}_{(j,k)}{\bf 1}\|_{1}\right).$
Summing up 9.6, Eq. 52 and Eq. 53 concludes the proof.
Supplementary material
In this supplementary material, we present a concise framework for MOT problems with global entropy regularization. Extending our previous work [fan2021complexity], this framework allows for marginal constraints on the joint distributions. Furthermore, it incorporates the capacity for cost functions to depend on multiple marginals, a feature previously implied but now explicitly defined and integrated into our algorithm.
A line of recent works have drawn a connection between entropy-regularized multi-marginal optimal transport problems (4) with a graph-structured cost (8) and probabilistic graphical models (PGMs) [HaaSinZha20, fan2021complexity, HaaRinChe20, SinHaaZha20, ringh2023mean, zhou2022efficient]. In this section we review and generalize these connections and the computational methods for solving these types of optimal transport problems.
10.1 A PGM formulation for multi-marginal optimal transport
A PGM provides a compact representation for the joint distribution of a collection of random vectors, which have dependencies that can be described by a graph [KolFri09]. Define the tensors $K_{\alpha}({\bf x}_{\alpha})=\exp(-C_{\alpha}({\bf x}_{\alpha})/\epsilon)$ for $\alpha\in F$, where $C_{\alpha}$ are the cost tensors in Eq. 8, and $\epsilon$ is the regularization parameter in Eq. 4. Then, $\bf K$} in Eq. 6 has the form ${\bf K}({\bf x})=\prod_{\alpha\in F}K_{\alpha}({\bf x}_{\alpha})$. Moreover, to define the entropy regularization Eq. 5 for the graph-structured problem Eq. 9 we define the tensor $\bf M$} as ${\bf M}({\bf x})=\prod_{\alpha\in F_{\Gamma}}\bm{\mu}_{\alpha}({\bf x}_{\alpha})$. The factorizations of the tensors $\bf K$} and $\bf M$} can be used to define the joint probability distribution of a PGM,
(54) $p({\bf x})=\frac{1}{Z}\left(\prod_{\alpha\in F}K_{\alpha}({\bf x}_{\alpha})\right)\left(\prod_{\alpha\in F_{\Gamma}}\bm{\mu}_{\alpha}({\bf x}_{\alpha})\right),$
where $Z$ is a normalization constant. Moreover, note that the solution tensor of (4) can be written as ${\mathbf{B}}={\bf K}\odot{\bf U}\odot{\bf M}$ with ${\bf U}({\bf x})=\prod_{\alpha\in F_{\Gamma}}{\mathbf{u}}_{\alpha}({\bf x}_{\alpha})$. Thus, the scaling factors ${\mathbf{u}}_{\alpha}$, for $\alpha\in F_{\Gamma}$ can be interpreted as local potentials of the modified graphical model ${\bf K}({\bf x}){\bf U}({\bf x}){\bf M}({\bf x})$ The Iterative Scaling algorithm 1 finds potentials ${\mathbf{u}}_{\alpha}$, for $\alpha\in F_{\Gamma}$, such that the graphical model defined by the tensor ${\mathbf{B}}={\bf K}\odot{\bf U}\odot{\bf M}$ satisfies all the constraints $P_{\alpha}({\mathbf{B}})=\bm{\mu}_{\alpha}$, for $\alpha\in F_{\Gamma}$.
10.2 Iterative Scaling Belief Propagation
Each step in the Iterative Scaling algorithm involves calculating a projection $P_{j}({\bf K}\odot{\bf U}\odot{\bf M})$, which in terms of the PGM means inferring the distribution of the $j$-th variable. This Bayesian inference problem can be addressed by the Belief Propagation algorithm [mezard2009information, Part D], [KolFri09, Sec. 20.5.1]. Belief propagation is known to converge when the underlying graph is a tree, i.e., does not have any loops, and reduces the complexity of computing the projections in the Iterative Scaling algorithm immensely. More precisely, whereas a brute force approach to compute the projections Eq. 3 scales exponentially, Belief Propagation scales only linearly in the number of marginals [HaaSinZha20].
In [fan2021complexity], a similar approach has be utilized for general graphical models with possible loops, by partitioning the graph as a so-called junction tree. A junction tree (also called tree decomposition) describes a partitioning of a graph, where several nodes are clustered together, such that the interactions between the clusters can be described by a tree structure.
Definition 10.1 (Junction tree [dawid1992applications]).
(55) $\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle+\epsilon{\mathcal{H}}({\mathbf{B}}{|{\bf M}})\qquad\text{ subject to }P_{c}({\mathbf{B}})=\bm{\mu}_{c},\text{ for }c\in\mathcal{C}_{\Gamma},$
(56) ${\bf C}({\bf x})=\sum_{c\in\mathcal{C}}{\bf C}_{c}({\bf x}_{c}).$
Example 10.2.
Some examples of valid junction trees for our method are shown in Fig. 8 (for the Euler flow example in Fig. 1), Fig. 9 (for the Wasserstein least squares problem Eq. 15 in Fig. 1) , and Fig. 10 (for the Wasserstein splines problem in Fig. 1).
1,2,31,2,31,31,31,3,41,3,41,41,41,4,51,4,51,J1,J1,J-1,J1,J1,J1,J1,J222233334444J1J-1J1J-1
Figure 9: Junction tree for the graph for the Wasserstein least square problem from Fig. 1. The junction tree is constructed such that the leaves of the tree correspond to the constraints in the least square problem Eq. 15.
[…]
Figure 10: Junction tree of Wasserstein splines in Fig. 1.
while not converged do
  -0.3cm
(57) $m_{i\to j}(s_{ij})$ $\leftarrow\sum_{x_{c_{i}}\setminus s_{ij}}{\bf K}_{c_{i}}({\bf x}_{c_{i}})\prod_{\ell\in N(i)\setminus j}m_{\ell\to i}(s_{\ell i}),\qquad c_{i}\in{\mathcal{C}}\setminus{\mathcal{C}}_{\Gamma}$
(58) $m_{i\to j}(s_{ij})$ $\leftarrow\mu_{c_{i}}({\bf x}_{c_{i}})/m_{j\to i}({\bf x}_{c_{i}}),\qquad c_{i}\in{\mathcal{C}}_{\Gamma}.$
return $m_{i\to j}(s_{ij}),\ m_{j\to i}(s_{ji}),\ \forall(i,j)\in\mathcal{E}$
Similarly to the factorization Eq. 54, we can then write the probability distribution over the PGM following the junction tree decomposition $\mathcal{T}=(\mathcal{C},\mathcal{E})$ as a product $p({\bf x})=\frac{1}{Z}\left(\prod_{c\in\mathcal{C}}{\bf K}_{c}({\bf x}_{c})\right)\left(\prod_{c\in\mathcal{C}_{\Gamma}}\bm{\mu}_{c}({\bf x}_{c})\right).$ Let $N(i)$ be the neighboring nodes of the node $i$. By extending the Iterative Scaling Belief Propagation methods in [HaaSinZha20, fan2021complexity], we propose Algorithm 4 to solve Eq. 55. Upon convergence of Algorithm 4, the optimal transport plans on the clusters are given by
(59) ${\mathbf{B}}_{c}({\bf x}_{c})={\bf K}_{c}({\bf x}_{c})\prod_{\ell\in N(i)}m_{\ell\to i}(s_{\ell i}).$
The algorithm sends messages between neighboring cliques. The messages are thus of the same size as the separation sets. Recall that the leaves of the junction tree are associated with constraints in the optimal transport problem. Thus, the messages sent from these leaves correspond to an update step similar to the one in Algorithm 1. The messages sent within the graph encode the computation of the projections. More precisely, picking one leaf $c$ and sending the messages from all other leaves across the separation sets towards our leaf amounts to one exact computation of the projection $P_{c}({\bf K}\odot{\bf U}\odot{\bf M})$. However, the algorithm converges with any message schedule [HaaSinZha20, fan2021complexity].
Note that in the case where all cost are pair-wise and constraints act on one-marginals of the transport tensors, Algorithm 4 boils down to the Iterative Scaling Belief Propagation algorithms in [HaaSinZha20, fan2021complexity]. Here, we have extend on these methods and provide a more general algorithm for the case where we may have constraints on joint distributions, and where cost terms can depend on more than two marginals.
The Iterative Scaling Belief Propagation algorithm utilizes the graphical structure in problem Eq. 55-Eq. 56 to significantly improve on the complexity of the Sinkhorn iterations. Recall that the computational bottleneck of Algorithm 1 lies in the computation of the projections, which require $\mathcal{O}(d^{J})$ operations. Algorithm 4 reduces this complexity by utilizing the graphical structure for computing the projections. The complexity of the problem depends then on the size of the messages, i.e., the size of the separation sets. This quantity can also be described by the so-called tree-width.
Definition 10.3 (Tree width [halin1976s]).
In order for Algorithm 4 to be efficient we thus would like to pick a junction tree factorization with small tree-width. Given a graph, a junction tree decomposition that minimizes the tree-width is also called a minimal junction tree. In general, finding a junction tree with minimal width is a NP hard problem [KolFri09]. However, for all applications in this paper it is easy to find minimal junction trees.
Example 10.4.
The junction trees in Figures 8, 9, and 10 are minimal junction trees for the respective graphs. Since the cliques in the former two junction trees are of cardinality 3, the respective graphs’ tree-widths are 2. Similarly, the last graph’s tree width is 3.
For Algorithm 4, we can directly apply the complexity analysis in [fan2021complexity]. We will discuss the complexity result in Remark 3.12.